Please show steps
EXERCICE 2 Convert to binary (2s complement) using a compact notation (minimum number of digits). Number in base 10 Number in base 2 (2s complement) +126.5 -25.8125 1.375 +10.37890625 13.62109375 15.61328125 2.99609375 EXERCICE 3 Give the result of the following set of additions in 8-bit 2s complement. Addends are also in 8-bit 2s complement. Indicate by YES or NO if an overflow occurs. Addition Result Overflow ? 0011 1000 0110 0000 1011 1000 1110 0000 1100 1000 1100 1000 1000 1000 1000 1000
Answer #1

1. 125 Fir st we neel to Convert ti26 5 into binamy no srsiem to refe sent )sim ve use o au in Gae o )weuses 2 1 26 2 6 3 o 2 31 1 2. 3 1 ve an conite +26.5 111 10 1 ( Binry) 2.25 8125 2 25 2 21 25 in binar 1 1001 0.625 ×2=1.25 o25 X2 0.50 o 0. 5 X2 1.0 1 2.6 Commfe mend ot (mool. 110): 000110.0011 (nn) .5 14: + 10-3-8906 2 0. 7573 i 25 x2『 . 51 56 25 0, 5, 5625 x2· 1-03125 0-03192-0 0623o 00625 χ2 0.125 o s .. coe Crnconite t10-3189062?→01010. 01100001 , . 2 s comple nod o-Q1010 01100001)ブ10 101 . I oo)1111 (in binary) 01010 01100001)プ10101. 100)1111 肋S ) 5.-13 6216 937S 2/3 O. 621093-5x2-J-243875 an à 0.24 3175 χ2=o-48孖s 0 951 0 902 メ2 e l、PO4 o 104 o 60g o 216 0 432 0864 0.798 x21216 x2 0-932 x20 264 0 0 936 0 912 0 324 0-648 21819 x2 698 in this case ae enve upto host. 8 bit.ne nd 아(11101-1001,川)ナ00010-01100001 ( b -15.G132 8125 22656 2-0-453125o 0 453121 90625o 0.90625 Y2:1.9125 0 8125 0.62S 0-25 1. 1 γ2-1.625 gcorn nement Coan Plement of (111),. )6011101) →ooooo : O11 oool, (And) C111), 18011101/>Oooooー01100011 (A4) 2./2 0.914375 x2:1 の6375 0 937S O 875 x2- ):75 co e com a0nite -9.ggCo g 3 75 →)10.111) 111 1 (inbinann.t : . 2s Compleme d δ1(110. 1111}111)→ 001.00000001 (Arg)001 o o 0 1 1 2 Comflemend ot h Result 11100000 over fiow YES 2 0 I1I 1161 11 10 000 3. O0I1 10oo 011o 0000 overflow 4 1011 1000 110 0oo0 om p Over frow YES s. I111 1000 2 icom femenk 아 he resu+-00101 ooo ovenfrow - YES 00 1000 11O 010000 2 sCo m ple ment ot te resott- 10000 . ovem flou - yes 1000 1000 l00o 1000 2s Compie ment ot he resolt- 1110000 over flow YEs

For the first part first we need to find the binaar value of the given no and then convert that into his 2's complement from.

And for the second part simply we need to add the numbers and if no of bit>8

Then overflow occurr else not.And then convert the results into 8 bit 2's complement representation.